# Learning Fairer Representations with FairVIC

### v2

@@ 5 Experiments
[…]
@@ 5.1 Datasets
We evaluate FairVIC on five datasets that are used as benchmarks in bias mitigation evaluation due to their known biases towards certain subgroups of people within their sample population. These datasets allow for highlighting the generalisable capabilities of FairVIC across different demographic disparities.
  • Tabular datasets. The main body of evaluation is done using three tabular datasets: Adult Income (Becker and Kohavi, 1996), COMPAS (Angwin et al., 2022), and German Credit (Hofmann, 1994), all of which are binary classification tasks. Adult Income aims to predict whether an individual’s income is over \$$\$50$K or not. It is known for its gender and racial biases in economic disparity. The Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) dataset is frequently used for evaluating debiasing techniques. It has a classification goal of predicting recidivism risks and is infamous for its racial biases. Finally, the German Credit dataset was used to assess creditworthiness by classifying individuals as either good or bad credit risks, with known biases related to age and gender (Kamiran and Calders, 2009).
  • Language datasets. To show the ability of FairVIC to work for different data modalities, we also utilise CivilComments-WILDS (Koh et al., 2021) and BiasBios (De-Arteaga et al., 2019) – natural language datasets. We collected a stratified sample of 50,000 text instances from each dataset, ensuring equal representation of both binary classification outcomes. CivilComments-WILDS is comprised of a collection of comments on the Civil Comments platform, with a binary classification goal to label each comment as toxic or non-toxic. We take ethnicity as the protected characteristic where comments are marked as white or non-white. BiasBios is a collection of professional biographies, with gender as the protected characteristic. We take a set of privileged and unprivileged professions, typically associated with a certain gender, as our binary classification goal.
@@ 5.2 Comparable Techniques
  • Adversarial Debiasing. This method leverages an adversarial network that aims to predict protected characteristics based on the predictions of the main model. The primary model seeks to maximise its own prediction accuracy while minimising the adversary’s prediction accuracy (Zhang et al., 2018).
  • Exponentiated Gradient Reduction. This technique reduces fair classification to a sequence of cost-sensitive classification problems, returning a randomised classifier with the lowest empirical error subject to a chosen fairness constraint (Agarwal et al., 2018).
  • Meta Fair Classifier. This classifier allows a fairness metric as an input and optimises the model with respect to regular performance and the chosen fairness metric (Celis et al., 2019).
  • Fair MixUp. This technique generates synthetic samples by linearly interpolating between pairs of training data points by protected attribute to smooth decision boundaries. The loss function is then further constrained by a fairness metric (Chuang and Mroueh, 2021).
  • FairGBM. This method uses a gradient-boosting decision tree model that integrates fairness constraints directly into the boosting process by adjusting the loss function to account for fairness metrics (Cruz et al., 2023).
Alongside these comparisons, a baseline neural network model using only binary cross-entropy loss was implemented, which exhibits the biases present in the datasets used. Details on the neural network architecture/ hyperparameters used for both the baseline model and the FairVIC model can be found in Appendix A.2.
@@ 6.1 Core Results Analysis
Table 1: FairVIC accuracy and fairness results, compared with the baseline model, and five other comparable methods for bias mitigation in-processing for each of the three tabular datasets.
@@
German Credit FairVIC     0.7250 $\pm$ 0.0239 0.8108 $\pm$ 0.0237     0.1443 $\pm$ 0.0796 0.1017 $\pm$ 0.0464 0.0016 $\pm$ 0.0604 1.0037 $\pm$ 0.0764
To assess the prediction and fairness performance of FairVIC [^2] FairVIC weights $L_{\text{acc}}=0.2$, $L_{\text{var, inv}}=0.1$, and $L_{\text{cov}}=0.6$ for the Adult Income dataset, and $L_{\text{acc, var, inv}}=0.1$, and $L_{\text{cov}}=0.7$ for the COMPAS and German Credit datasets, see Appendix B.3 for discussion on these selections. and state-of-the-art approaches, we test all methods across each tabular dataset to enable a fair comparison. Table 1 shows these results. We have also provided Figure 1, which visualises the absolute difference from the ideal value of each metric, highlighting how far each method deviates from perfect accuracy and fairness on each tabular dataset.
@@
Across all three datasets, the baseline performs poorly in fairness but obtains higher performance scores, which is expected. For example, in the Adult Income dataset, the baseline model shows a relatively high accuracy ($0.8444$), while exhibiting poor fairness with regard to Disparate Impact ($0.2853$). The baseline highlights the need for a bias mitigation approach that works across all metrics simultaneously, as the lower bias in terms of Equalised Odds ($0.1330$) and Absolute Odds ($0.1172$) alone could misleadingly suggest that the model is fair, when in reality, the bias may only become evident when captured through a different perspective. This means that approaches relying on a single fairness constraint, such as Exponentiated Gradient Reduction, often fail to address significant bias present in the data.
@@
Overall, FairVIC outperforms all other comparable methods by demonstrating consistent improvements in both fairness and accuracy retention. As seen in Figure 1(a), our FairVIC model achieves the lowest cumulative absolute error from perfect accuracy and fairness in the Adult Income dataset, effectively balancing the fairness-accuracy trade-off. The trend is also consistent across the COMPAS and German Credit datasets as seen in Figures 1(b) and 1(c). This again exemplifies the ability of FairVIC’s approach to generalise across datasets, making it a more versatile solution.
(a) Adult Income dataset.
(b) COMPAS dataset.
(c) German Credit dataset.
@@
Figure 1: Absolute differences from the ideal value (e.g., perfect accuracy and fairness) in performance (left) and fairness (right) metrics of comparable techniques, sorted in ascending order on all three tabular datasets.
@@
Other comparable methods are generally not as effective as FairVIC, each exhibiting different shortcomings. For instance, MetaFair often struggles to improve even upon the baseline in cumulative absolute difference from the ideal value, and many techniques struggle to balance the improvements across all fairness metrics, often prioritising Equalised and Absolute Odds over Disparate Impact, particularly in the Adult Income dataset. Similarly, FairMixUp, though initially promising and achieving second place after FairVIC in the COMPAS and German Credit datasets, fails to maintain its performance on the Adult Income dataset, where its results only just beat the baseline. In many cases, such as FairMixUp on the COMPAS and German Credit datasets, comparable techniques improve fairness but at the cost of accuracy, failing to achieve a balanced tradeoff.
Overall, FairVIC’s ability to consistently balance the trade-off between fairness and accuracy, adapt to various datasets, and handle all fairness metrics comprehensively makes it the most effective method. Its consistent performance across different datasets, as evidenced by the lowest cumulative absolute error in performance and fairness, solidifies its superiority over other comparable methods.
@@ 6.2 Individual Fairness Analysis
To emphasise further FairVIC’s ability to perform well across all fairness metrics, we also evaluate upon individual fairness by outputting the results of the counterfactual model, as described in Section 3.3. The full results, alongside the absolute difference in averages for each metric across the regular and counterfactual models, are seen in Table 5 in Appendix B.2.
The FairVIC model shows considerable promise in enhancing individual fairness across different datasets when compared with the baseline models. The counterfactual results from the FairVIC model with invariance term weighted heavily (FairVIC Invariance) exhibits lower absolute differences in metrics across all datasets. For example, in the German Credit dataset, the mean absolute difference across all six metrics between the regular and the counterfactual baseline model is $\approx 0.0277$, while for FairVIC Invariance’s regular and counterfactual models it is lower at $\approx 0.0108$. This suggests a more stable and fair performance under counterfactual conditions. This capability highlights FairVIC’s strength in not only addressing group fairness but also ensuring that individual decisions remain consistent and fair when hypothetical scenarios are considered. In the FairVIC model with recommended lambdas, we prioritise group fairness so invariance is weighted less. Even with this lower invariance weighting, FairVIC still achieved improved individual fairness.
Table 2: FairVIC and baseline comparison results of both performance and fairness for the CivilComments-WILDS and BiasBios datasets, including FairVIC’s counterfactual (CF) model results and the absolute differences (ADs) between each model.
@@
CivilComments-WILDS FairVIC AD     0.0921 $\pm$ 0.0907 0.0892 $\pm$ 0.2381     0.0141 $\pm$ 0.0711 0.0077 $\pm$ 0.0751 0.0329 $\pm$ 0.0648 0.0657 $\pm$ 0.1844
BiasBios Baseline     0.8818 $\pm$ 0.0034 0.8811 $\pm$ 0.0044     0.0558 $\pm$ 0.0103 0.0456 $\pm$ 0.0083 -0.2489 $\pm$ 0.0098 0.6038 $\pm$ 0.0159
Baesline CF     0.8794 $\pm$ 0.0041 0.8797 $\pm$ 0.0032     0.0563 $\pm$ 0.0153 0.0461 $\pm$ 0.0120 0.2481 $\pm$ 0.0119 1.6401 $\pm$ 0.0292
@@
FairVIC     0.8653 $\pm$ 0.0070 0.8646 $\pm$ 0.0059     0.0992 $\pm$ 0.0258 0.0830 $\pm$ 0.0147 -0.1217 $\pm$ 0.0158 0.7817 $\pm$ 0.0352
FairVIC CF     0.8587 $\pm$ 0.0071 0.8580 $\pm$ 0.0074     0.1472 $\pm$ 0.0304 0.1193 $\pm$ 0.0186 0.0844 $\pm$ 0.0194 1.1890 $\pm$ 0.0568
FairVIC AD     0.0066 ± 0.0066 0.0066 ± 0.0062     0.0480 ± 0.0392 0.0363 ± 0.0193 0.2061 ± 0.0276 0.4074 ± 0.0735
@@ 6.3 Lambda Ablation Study Analysis
The FairVIC loss terms are combined with binary cross entropy for training the neural network to enable optimisation of both accuracy and fairness, minimising the trade-off. The effect of FairVIC on the overall loss function can be increased and decreased by changing the weight $\lambda$ for each FairVIC term. To evaluate this effect, we train a number of neural networks with the architecture described in Appendix A.2, with a different $\lambda_{\text{acc}}$ weighting each time. In this experiment, we evaluate the effect of weighting the FairVIC loss terms equally, so that $\lambda_{\text{var}}=\lambda_{\text{inv}}=\lambda_{\text{cov}}=\frac{(1-\lambda_{\text{acc}})}{3}$, where $10<\lambda_{\text{acc}}<$ 1. The performance and fairness measures for each model are listed in Table 6 in Appendix C, and visualisations for the absolute difference in performance and fairness from ideal values for each run are visualised in Figure 6 in Appendix C.
@@
To evaluate upon the effect of each individual VIC term within the loss function, we can suppress the lambda terms from two out of three of variance, invariance, and covariance to leave only one remaining. We keep $\lambda_{\text{acc}}=0.1$ since the previous lambda experiment showed this to be most effective and revealing in terms of the effect on fairness, while the chosen FairVIC loss term is assigned a weighting of $0.9$. Similarly, we can also suppress a single term at a time, assigning two out of the three VIC terms a weighting of $0.45$. The performance and fairness results for each experiment with different weightings are listed in Table 7.
@@
6.4 Language Dataset Results
To show FairVIC’s versatility across data modalities, our approach was applied to the CivilComments-WILDS and BiasBios datasets. The results are shown in Table 2, where FairVIC uses the lambdas $L_{\text{acc, var, inv}}=0.1$, and $L_{\text{cov}}=0.7$ for the CivilComments-WILDS dataset and $L_{\text{acc, var, inv}}=0.05$, and $L_{\text{cov}}=0.85$ for the BiasBios dataset.
From Table 2, the same trend can be seen as in the tabular dataset results, where FairVIC gives fairer results across all tested fairness metrics, the most notable being the improvement to disparate impact from $1.9390$ to $1.1344$ in the CivilComments-WILDS and $0.6038$ to $0.7817$ in the BiasBios datasets. In terms of individual fairness, the CivilComments-WILDS’ baseline model has a mean absolute difference across every metric between the regular and counterfactual model of $\approx 0.3303$ and $\approx 0.0503$ for FairVIC. For the BiasBios dataset, the baseline model has a mean absolute difference across every metric between the regular and counterfactual model of $\approx 0.2563$ and $\approx 0.1185$ for our FairVIC model. Therefore, FairVIC is not confined to one modality, due to its ability to effectively reduce both individual and group fairness, while not seeing a drop of more than $3.81\%$ and $1.87\%$ in accuracy for the CivilComments-WILDS and BiasBios datasets respectively. The use of a different model architecture also proves FairVIC’s adaptability to be utilised within different neural networks.

### v3

@@ 5 Experiments
In our experimental evaluation, we assess the performance of FairVIC [^1] The code for our FairVIC implementation is provided at: https://github.com/CharmaineBarker/FairVIC. against a set of comparable in-processing bias mitigation methods on a series of datasets known for their bias. Here, we describe the datasets used and the methods we compare against.
@@ 5.1 Datasets
We evaluate FairVIC on seven widely used bias mitigation benchmarks across tabular, text, and image modalities. These datasets contain known demographic disparities, allowing us to assess FairVIC’s generalisability. For each, we designate one attribute as the protected characteristic on which fairness is to be improved.
  • Tabular datasets. We use Adult Income [33], COMPAS [34], and German Credit [35], all binary classification tasks with known biases. Adult Income predicts whether income exceeds $\mathdollar 50$K and exhibits gender and racial bias. COMPAS predicts recidivism risk and is notorious for racial bias. German Credit classifies creditworthiness, with biases related to age and gender [36].
  • Language datasets. CivilComments-WILDS [37] and BiasBios [38] are used to assess FairVIC on text data. We sample 50,000 stratified examples per dataset to ensure balance. CivilComments classifies online comments as toxic or non-toxic, using race as the protected attribute. BiasBios consists of professional biographies classified into favourable vs. unfavourable occupations (e.g., surgeon vs. nurse), with gender as the protected feature.
  • Image datasets. CelebA [39] contains celebrity portraits labelled for attributes such as blond hair; we predict whether a person has blond hair or not, using gender as the protected attribute. UTKFace [40] includes facial images labelled with age and race; we predict whether a subject is above or below 30, with race as the protected attribute.
@@ 5.2 Comparable Techniques
  • Adversarial Debiasing. This method leverages an adversarial network that aims to predict protected characteristics based on the predictions of the main model. The primary model seeks to maximise its own prediction accuracy while minimising the adversary’s prediction accuracy [19].
  • Exponentiated Gradient Reduction. This technique reduces fair classification to a sequence of cost-sensitive classification problems, returning a randomised classifier with the lowest empirical error subject to a chosen fairness constraint [18].
  • Meta Fair Classifier. This classifier allows a fairness metric as an input and optimises the model with respect to regular performance and the chosen fairness metric [17].
  • Fair MixUp. This technique generates synthetic samples by linearly interpolating between pairs of training data points by protected attribute to smooth decision boundaries. The loss function is then further constrained by a fairness metric [23].
  • FairGBM. This method uses a gradient-boosting decision tree model that integrates fairness constraints directly into the boosting process by adjusting the loss function to account for fairness metrics [24].
A baseline neural network trained with binary cross-entropy was also implemented to reflect dataset biases. Model architecture and hyperparameters for both baseline and FairVIC are detailed in Appendix A.3.
@@ 6.1 Core Results Analysis
To assess the prediction and fairness performance of FairVIC [^2] See Table 6, Appendix B.3 for our FairVIC lambdas, and discussion on these selections. and state-of-the-art approaches, we test all methods across each tabular dataset to enable a fair comparison. Table 1 shows these results. We have also provided Figure 3, which visualises the absolute difference from the ideal value of each metric, highlighting how far each method deviates from perfect accuracy and fairness on each tabular dataset.
Table 1: Accuracy and fairness results of FairVIC versus the baseline and five in-processing methods on three tabular datasets. Bold indicates the best overall trade-off between fairness and accuracy, reflecting our aim of holistic improvement (see Figure 3). Underlined values show the best score for each individual metric.
Performance Metrics Fairness Metrics
@@
German
Credit
@@
Across all three datasets, the baseline performs poorly in fairness but obtains higher performance scores, which is expected. For example, in the Adult Income dataset, the baseline model shows a relatively high accuracy ($0.8444$), while exhibiting poor fairness with regard to Disparate Impact ($0.2853$). The baseline highlights the need for a bias mitigation approach that works across all metrics simultaneously, as the lower bias in terms of Equalised Odds ($0.1330$) and Absolute Odds ($0.1172$) alone could misleadingly suggest that the model is fair, when in reality, the bias may only become evident when captured through a different perspective. Single-metric approaches like EGR often fail to address significant bias.
@@
Overall, FairVIC outperforms all other comparable methods by demonstrating consistent improvements in both fairness and accuracy retention. As seen in Figure 1, our FairVIC model achieves the lowest cumulative absolute error from perfect accuracy and fairness in the Adult Income dataset, effectively balancing the fairness-accuracy trade-off. The trend is also consistent across the COMPAS and German Credit datasets seen in Figure 3, Appendix B. This further highlights FairVIC’s ability to generalise across datasets.
@@
Figure 1: A qualitative analysis of the absolute differences from the ideal value (e.g., perfect accuracy and fairness) in performance (left) and fairness (right) metrics of comparable techniques on the Adult Income dataset.
@@
Other comparable methods are generally not as effective as FairVIC, each exhibiting different shortcomings. For instance, MetaFair often struggles to improve even upon the baseline in cumulative absolute difference from the ideal value, and many techniques struggle to balance the improvements across all fairness metrics, often prioritising Equalised and Absolute Odds over Disparate Impact, particularly in the Adult Income dataset. Similarly, FairMixUp, though initially promising and achieving second place after FairVIC in the COMPAS and German Credit datasets, fails to maintain its performance on the Adult Income dataset, where its results only just beat the baseline. In many cases, such as FairMixUp on the COMPAS and German Credit datasets, comparable techniques improve fairness but at the cost of accuracy, failing to achieve a balanced tradeoff. Exponentiated Gradient Reduction (EGR) sees similar struggles. While performing the best in terms of each individual fairness metric in the COMPAS dataset, it does this at the expense of accuracy, where it sees a drop of $15.83\%$ in accuracy. Further to this, EGR then performs poorly on the Adult Income dataset, only achieving a disparate impact of 0.4602, suggesting inability to generalise across datasets.
Following this paper’s objective to create an approach that performs well across multiple fairness metrics without significantly compromising accuracy, we find that FairVIC demonstrates a consistent ability to balance fairness and accuracy across diverse datasets. Its adaptability, strong performance on all fairness metrics, and robustness to dataset shifts position it as the most effective method overall. This is further supported by its consistently low cumulative absolute error across both fairness and performance measures, highlighting its advantage over existing in-processing techniques.
@@ 6.2 Individual Fairness Analysis
To emphasise further FairVIC’s ability to perform well across all fairness metrics, we also evaluate upon individual fairness by outputting the results of the counterfactual model, as described in Section 3.3. The full results, alongside the absolute difference in averages for each metric across the regular and counterfactual models, are seen in Table 5, Appendix B.2.
The FairVIC model shows considerable promise in enhancing individual fairness across different datasets when compared with the baseline models. The counterfactual results from the FairVIC model with invariance term weighted heavily (FairVIC Invariance) exhibits lower absolute differences in metrics across all datasets. For example, in the German Credit dataset, the mean absolute difference across all six metrics between the regular and the counterfactual baseline model is $0.0277$, while for FairVIC Invariance’s regular and counterfactual models it is lower at $0.0108$. This suggests a more stable and fair performance under counterfactual conditions. This capability highlights FairVIC’s strength in not only addressing group fairness but also ensuring that individual decisions remain consistent and fair when hypothetical scenarios are considered. In the FairVIC model with recommended lambdas, we prioritise group fairness so invariance is weighted less. Even with this lower invariance weighting, FairVIC still achieved improved individual fairness.
Table 2: FairVIC and baseline comparison results of both performance and fairness for each of the four text and image datasets, including FairVIC’s counterfactual (CF) model results and the absolute differences (ADs) between each model.
Performance Metrics Fairness Metrics
@@
CivilComments
-WILDS
FairVIC AD 0.0921 $\pm$ 0.0907 0.0892 $\pm$ 0.2381 0.0141 $\pm$ 0.0711 0.0077 $\pm$ 0.0751 0.0329 $\pm$ 0.0648 0.0657 $\pm$ 0.1844
BiasBios
Baseline 0.8818 $\pm$ 0.0034 0.8811 $\pm$ 0.0044 0.0558 $\pm$ 0.0103 0.0456 $\pm$ 0.0083 -0.2489 $\pm$ 0.0098 0.6038 $\pm$ 0.0159
Baseline CF 0.8794 $\pm$ 0.0041 0.8797 $\pm$ 0.0032 0.0563 $\pm$ 0.0153 0.0461 $\pm$ 0.0120 0.2481 $\pm$ 0.0119 1.6401 $\pm$ 0.0292
@@
FairVIC 0.8668 $\pm$ 0.0070 0.8633 $\pm$ 0.0087 0.0970 $\pm$ 0.0167 0.0830 $\pm$ 0.0084 -0.1220 $\pm$ 0.0110 0.7744 $\pm$ 0.0207
FairVIC CF 0.8601 $\pm$ 0.0069 0.8582 $\pm$ 0.0059 0.1464 $\pm$ 0.0222 0.1181 $\pm$ 0.0112 0.0862 $\pm$ 0.0117 1.1945 $\pm$ 0.0304
FairVIC AD 0.0066 $\pm$ 0.0063 0.0051 $\pm$ 0.0070 0.0493 $\pm$ 0.0240 0.0350 $\pm$ 0.0146 0.2082 $\pm$ 0.0132 0.4201 $\pm$ 0.0295
Baseline 0.8765 $\pm$ 0.0048 0.8776 $\pm$ 0.0057 0.2019 $\pm$ 0.0516 0.1268 $\pm$ 0.0249 -0.4672 $\pm$ 0.0124 0.4353 $\pm$ 0.0208
Baseline CF 0.8747 $\pm$ 0.0063 0.8724 $\pm$ 0.0106 0.2286 $\pm$ 0.0495 0.1396 $\pm$ 0.0232 0.4676 $\pm$ 0.0096 2.2250 $\pm$ 0.1288
Baseline AD 0.0018 $\pm$ 0.0055 0.0052 $\pm$ 0.0073 0.0267 $\pm$ 0.0510 0.0127 $\pm$ 0.0246 0.9348 $\pm$ 0.0176 1.7897 $\pm$ 0.1319
FairVIC 0.7571 $\pm$ 0.0286 0.7811 $\pm$ 0.0249 0.2788 $\pm$ 0.0869 0.1543 $\pm$ 0.0349 -0.1018 $\pm$ 0.0815 0.8259 $\pm$ 0.1596
FairVIC CF 0.7442 $\pm$ 0.0140 0.7804 $\pm$ 0.0180 0.3189 $\pm$ 0.0372 0.1686 $\pm$ 0.0148 0.0612 $\pm$ 0.0318 1.1956 $\pm$ 0.0909
CelebA
FairVIC AD 0.0129 $\pm$ 0.0142 0.0006 $\pm$ 0.0239 0.0401 $\pm$ 0.0584 0.0142 $\pm$ 0.0214 0.1630 $\pm$ 0.0958 0.3697 $\pm$ 0.1949
UTKFace
Baseline 0.6827 $\pm$ 0.0120 0.6839 $\pm$ 0.0225 0.0887 $\pm$ 0.0322 0.0631 $\pm$ 0.0260 0.1261 $\pm$ 0.0265 1.3136 $\pm$ 0.1024
Baseline CF 0.6723 $\pm$ 0.0230 0.6655 $\pm$ 0.0767 0.0988 $\pm$ 0.0218 0.0741 $\pm$ 0.0224 -0.1328 $\pm$ 0.0233 0.7416 $\pm$ 0.0868
Baseline AD 0.0104 $\pm$ 0.0198 0.0184 $\pm$ 0.0521 0.0100 $\pm$ 0.0106 0.0110 $\pm$ 0.0101 0.2589 $\pm$ 0.0450 0.5720 $\pm$ 0.1806
FairVIC 0.6308 $\pm$ 0.0288 0.5901 $\pm$ 0.1011 0.0647 $\pm$ 0.0335 0.0509 $\pm$ 0.0313 -0.0035 $\pm$ 0.0332 1.0076 $\pm$ 0.0770
FairVIC CF 0.6231 $\pm$ 0.0247 0.5713 $\pm$ 0.1176 0.0796 $\pm$ 0.0367 0.0645 $\pm$ 0.0325 0.0146 $\pm$ 0.0385 1.0097 $\pm$ 0.1332
FairVIC AD 0.0077 $\pm$ 0.0250 0.0188 $\pm$ 0.1106 0.0149 $\pm$ 0.0317 0.0136 $\pm$ 0.0251 0.0181 $\pm$ 0.0334 0.0021 $\pm$ 0.0894
6.3 Language and Image Dataset Results
To demonstrate FairVIC’s versatility, we apply it to CivilComments-WILDS and BiasBios. Results are shown in Table 2; our selected lambda configurations are given in Table 6. Across all fairness metrics, FairVIC improves upon the baseline, consistent with trends seen in the tabular datasets. Notably, Disparate Impact improves from $1.9390$ to $1.1344$ (CivilComments-WILDS), $0.6038$ to $0.7817$ (BiasBios), $0.4353$ to $0.8259$ (CelebA), and $1.3136$ to $1.0076$ (UTKFace). For individual fairness, the mean absolute difference between regular and counterfactual models is reduced substantially: from $0.3303$ to $0.0503$ on CivilComments-WILDS, $0.2563$ to $0.1207$ on BiasBios, $0.4618$ to $0.1001$ on CelebA, and $0.1468$ to $0.0125$ on UTKFace. These results confirm FairVIC’s effectiveness across modalities and architectures, improving both group and individual fairness with minimal accuracy trade-off.
@@ 6.4 Lambda Ablation Study Analysis
Finally, to better understand the internal dynamics of FairVIC, we conduct an ablation study on the $\lambda$ weightings of FairVIC’s loss terms to analyse their individual contributions. The FairVIC loss terms are combined with binary cross entropy for training the neural network to enable optimisation of both accuracy and fairness, minimising the trade-off. The effect of FairVIC on the overall loss function can be increased and decreased by changing the weight $\lambda$ for each FairVIC term. To evaluate this effect, we train a number of neural networks with the architecture described in Appendix A.3, with a different $\lambda_{\text{acc}}$ weighting each time. In this experiment, we evaluate the effect of weighting the FairVIC loss terms equally, so that $\lambda_{\text{var}}=\lambda_{\text{inv}}=\lambda_{\text{cov}}=\frac{(1-\lambda_{\text{acc}})}{3}$, where $0<\lambda_{\text{acc}}<1$. The performance and fairness measures for each model are listed in Table 7, Appendix C, and visualisations for the absolute difference in performance and fairness from ideal values for each run are visualised in Figure 7, Appendix C.
@@
To evaluate the effect of each individual VIC term within the loss function, we can suppress the lambda terms from two out of three of variance, invariance, and covariance to leave only one remaining. We keep $\lambda_{\text{acc}}=0.1$ since the previous lambda experiment showed this to be most effective and revealing in terms of the effect on fairness, while the chosen FairVIC loss term is assigned a weighting of $0.9$. Similarly, we can also suppress a single term at a time, assigning two out of the three VIC terms a weighting of $0.45$. The performance and fairness results for each experiment with different weightings are listed in Table 8.
@@
Based on these findings, we can offer our recommendations for default lambda values to achieve effective results. We suggest beginning with a configuration of $\lambda_{\text{acc}}=0.1$, $\lambda_{\text{var}}=0.1$, $\lambda_{\text{inv}}=0.1$, and $\lambda_{\text{cov}}=0.7$, which consistently performs well across datasets. This configuration provides a strong default, with further tuning advised based on the relative importance of group fairness, individual fairness, or accuracy in a given application. A detailed discussion of dataset-specific configurations and practical guidance for adjusting these weights is provided in Appendix B.3.